

********************************************************************
UARS WEEKLY STATUS REPORT - 05 MARCH 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 54% of
each day's data amounting to nearly 466 minutes of solar observing time
per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, Mg II,
occultation, solar pointing offset and high resolution continuous
monitoring scans for periods of UARS real-time telemetry or onboard
tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1994189. MEPS data files between
1994190 and 1994219 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 21 February 2004 (Orbit 68097, GMT Day 052) through
27 February 2004 (Orbit 68202, GMT Day 058).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 054/1220 68134 FULLTIME OPS
SUSIM 054/1542 68136 POWER ON
HALOE 056/1356 68165 POWER ON
HALOE 056/1356 68165 POWER OFF
HALOE 056/1453 68166 POWER ON
A Forward-to-Reverse Yaw Around maneuver (number 127) was successfully
executed in Orbit 68106 beginning 052/14:45:41Z and completing at
052/15:22:51Z on 21 February 2004. All instruments were in their off
or standby mode (except for SOLSTICE in day-only). Due to the SA
position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 57.62 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 26.08 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 21.09 and 20.51 percent
respectively. The next (reverse-to-forward) YAW is scheduled for 29
March 2004 (DOY 089).
The beta angle ranged between the angles of 01.0 degrees to 25.0
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 60 degrees on 11 March 2004 (DOY 071).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.96 to 1.27 degrees C. Additional
battery performance data for each battery on 20 February and 27
February 2004 are provided in Appendix D.
The clock error ranged from 10.7 to -09.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
HALOE = 03
SUSIM = 02
SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continued testing the new version of the CMS code this week.
The testing is scheduled to be completed by 05 March 2004. A
tentative date of 09 March 2004 has been set for a mini-ORR to review
the results of testing and the plans to utilize the new 3-TDRS
functionality.
MEETINGS
The FOT held a meeting on 26 February 2004 to discuss the CMS code
problem and the regression test plan for the following weeks.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 29 March 2004 (DOY 089). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments:
Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 27 February 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
68103 052/1031 04-062 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
68105 052/1353 04-061 MACS YAW -FWD TO REV
68121 053/1501 04-061 MPS VT LEVEL 5 (NORMAL)
68121 053/1545 04-061 MPS BATSOC
68123 053/1835 04-063 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
68134 054/1220 04-064 SOLSTICE FULLTIME OPS
68136 054/1542 04-065 SUSIM POWER ON (BEGIN)
68151 055/1539 04-066 SUSIM POWER ON (COMPLETE)
68165 056/1356 04-067 HALOE POWER ON
68165 056/1356 04-068 HALOE POWER OFF
68166 056/1453 04-067 HALOE POWER ON
68168 056/1829 04-069 HALOE POWER ON (PART 2)
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 27 February 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 27 February 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
68103 052/1034 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
68123 053/1835 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 27 February 2004
20 February 2004 - GMT Day 051 Beta = 02.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-44.8 +5.14/+3.87 +28.4/-12.8 25.8 1.04 20.0
3 +11.2/-50.4 +3.87/+2.60 +27.2/-12.8 25.6 1.04 19.5
27 February 2004 - GMT Day 058 Beta = 25.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 +16.8/-072.8 +5.46/+4.50 +28.8/-14.4 25.6 1.03 20.5
3 +11.2/-112.0 +4.50/+3.23 +28.4/-14.4 25.4 1.02 19.9
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
68092 051/1638 LV5 -> LV6
68121 053/1501 LV6 -> LV5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 27 February 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-013 FOT 68165 056/1311 35212 Multiple Data Dropouts
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-011 FOT See List Feb 04 30832 Generic Late Acquisition
04-013 FOT 68165 056/1311 35212 Multiple Data Dropouts